OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = 4/3 x + 8 y = 4/3 x + 3 y = - 3/4 x + 8 y = - 3/4 x + 3

OpenStudy (nurali):

y=mx+b to be parallel, slopes must be equal rearrange 4x-3y=-12 3y=4x+12 y=4/3x +4 slope (m) is 4/3 sub in points it has to pass through (x=-3, y=4) (4)=4/3(-3)+b y intercept (b) is 8 y=4/3x+8

OpenStudy (anonymous):

thank you see to me it looks so easy but then I try and do it and I do something wrong ugh thank you again :)

OpenStudy (nurali):

My Pleasure.